IF THE PRICE OF ORANGE REDUCES BY 20%A MAN CAN BUY 48 ORANGES MORE FOR RS 240.FIND THE ORIGINAL & REDUCED PRICE OF ORANGE PER DOZEN?

Answer by mukhopadhyay(490)

Assume that the regular price per orange is Rs x;
Number of ornages that can be bought for Rs 240 is (240/x);
If the price per orange reduces by 20%, the reduced price per orange is 80% of x = .8x
Based on the reduced price, number of oranges that can be bought for Rs 240 would be (240/x + 48)
So, Reduced price per orange * Number of oranges based on reduced price = 240
=> .8x(240/x + 48) = 240
=> 192 + 38.4x = 240
=> 38.4x = 48
=> x = 48/38.4 = 1.25
The original price of each orange = Rs 1.25
Answer:
The original price of orange per dozen = Rs 1.25 * 12 = Rs 15
The reduced price per orange is 80% of Rs 15 = Rs 12